OpenStudy (anonymous):

How would you solve 10/8-5i ?

geerky42 (geerky42):

Multiply both numerator and denominator by \(8+5i\)

geerky42 (geerky42):

Like this:\[\dfrac{10}{8-5i}\cdot\dfrac{8+5i}{8+5i} ~~=~~\dfrac{10(8+5i)}{(8-5i)(8+5i)} \]

OpenStudy (anonymous):

thank you so much

geerky42 (geerky42):

Yep

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

i meant 80+50i/89 but, thanks